# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ II.5 Ingoing Eddington coordinates
$r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (II.57)
@@
$\frac{\partial r_{*}}{\partial r}=1+\frac{1}{2\kappa_{+}\left(r-r_{+}\right)}-\frac{1}{2\kappa_{-}\left(r-r_{-}\right)}=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}$ (II.58)
@@
$\frac{\partial t}{\partial v}=1,\quad\frac{\partial r}{\partial v}=0$ (II.61)
$\frac{\partial t}{\partial r}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1},\quad\frac{\partial r}{\partial r}=1$ (II.62)
@@
$g_{vv}=\frac{\partial t}{\partial v}\frac{\partial t}{\partial v}g_{tt}+\frac{\partial r}{\partial v}\frac{\partial r}{\partial v}g_{rr}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (II.63)
$g_{rr}$ $=\frac{\partial t}{\partial r}\frac{\partial t}{\partial r}g_{tt}+\frac{\partial r}{\partial r}\frac{\partial r}{\partial r}g_{rr}$ (II.64)
$=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (II.65)
$+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}=0$ (II.66)
@@
$g_{rv}$ $=\frac{\partial t}{\partial r}\frac{\partial t}{\partial v}g_{tt}+\frac{\partial r}{\partial r}\frac{\partial r}{\partial v}g_{rr}$ (II.67)
$=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)=1$ (II.68)
@@
$ds^{2}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)dv^{2}+2dr\;dv+r^{2}d\Omega^{2}$ (II.69)
II.6 RN in double null Eddington coordinates
@@
$\frac{\partial r_{*}}{\partial r}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)^{-1}$ (II.70)
@@
$g_{r_{*}r_{*}}=g_{rr}\frac{\partial r}{\partial r_{*}}\frac{\partial r}{\partial r_{*}}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)$ (II.71)
@@
$ds^{2}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)\left(-dt^{2}+dr_{*}^{2}\right)+r^{2}(r_{*})d\Omega^{2}$ (II.72)
@@
$du\cdot dv=dr_{*}^{2}-dt^{2}$ (II.79)
Substituting this in the metric Eq. (II.72), the Reissner–Nordström geometry in double null coordinates is found to take the form:
$ds^{2}=\left(1-\frac{2M}{r(u,v)}+\frac{Q^{2}}{r^{2}(u,v)}\right)du\;dv+r^{2}(u,v)d\Omega^{2}$ (II.80)
@@
$g_{uv}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (II.81)
@@
$\frac{u+v}{2}=r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (II.82)
@@
$\frac{u+v}{2}=r_{*}=const-\frac{1}{2\kappa_{-}}\log\left|r-r_{-}\right|$ (II.83)
@@
$r-r_{-}=const\cdot e^{-2\kappa_{-}r_{*}}=const\cdot e^{-\kappa_{-}\left(u+v\right)}$ (II.84)
The metric component $g_{uv}$ (Eq. (II.81)) around $r=r_{-}$ is (recalling also Eq. (II.1,II.56)):
$2\frac{e^{S}L_{0}}{r}$ $=-2g_{uv}$ (II.85)
$=-\frac{\left(r-r_{-}\right)\left(r-r_{+}\right)}{r^{2}}\approx-\frac{\left(r-r_{-}\right)\left(r_{-}-r_{+}\right)}{r_{-}^{2}}$ (II.86)
$=2\kappa_{-}\left(r-r_{-}\right)$ (II.87)
@@
$const\cdot 2\kappa_{-}e^{-\kappa_{-}\left(u+v\right)}=2\kappa_{-}\left(r-r_{-}\right)\approx 2\frac{e^{s}L_{0}}{r_{-}}$ (II.88)
@@
$S=-\kappa_{-}\left(u+v\right)+const$ (II.89)
@@
$S_{,u}=S_{,v}=-\kappa_{-}$ (II.90)
near the IH. Taking the derivative of Eq. (II.84) with respect to $r_{*},$we get near the IH:
$r_{,r_{*}}=const\cdot e^{-2\kappa_{-}r_{*}}$ (II.91)
@@
$R_{,u}=R_{,r}r_{,r_{*}}r_{*,u}=2r\cdot const\cdot e^{-2\kappa_{-}r_{*}}\frac{1}{2}=const\cdot e^{-2\kappa_{-}r_{*}}\rightarrow 0$ (II.92)
@@
$R_{,v}=R_{,r}r_{,r_{*}}r_{*,v}=const\cdot e^{-2\kappa_{-}r_{*}}\rightarrow 0$ (II.93)
@@
$r_{*}=const+\frac{1}{2\kappa_{+}}\log\left|r-r_{+}\right|$ (II.94)
@@
$r-r_{+}=const\cdot e^{2\kappa_{+}r_{*}}=const\cdot e^{\kappa_{+}\left(u+v\right)}$ (II.95)
@@
$2\frac{e^{S}L_{0}}{r}$ $=-2g_{uv}$ (II.96)
$=-\frac{\left(r-r_{+}\right)\left(r-r_{-}\right)}{r^{2}}\approx-\frac{\left(r-r_{+}\right)\left(r_{+}-r_{-}\right)}{r_{-}^{2}}$ (II.97)
$=-2\kappa_{+}\left(r-r_{+}\right)$ (II.98)
@@
$const\cdot 2\kappa_{+}e^{\kappa_{+}\left(u+v\right)}=2\kappa_{+}\left(r-r_{+}\right)\approx 2g_{uv}$ (II.99)
II.7 Vaidya metric
@@
$ds^{2}=-\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (II.100)
Note that this metric, when taking $M(v)=const$ and $Q(v)=const$, is exactly equal to Eq. (II.69) ingoing Eddington metric.
The $G_{vv}$ Einstein tensor of the Vaidya metric is:
$G_{vv}=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)+\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{Q^{2}}{r^{4}}$ (II.101)
The $G_{rv}$ Einstein tensor of the Vaidya metric is:
$G_{rv}=-\frac{Q^{2}}{r^{4}}$ (II.102)
We wish to know $G_{vv}$ in double null coordinates $(v,u)$, which we’ll mark $(V,U)$ for clarity, instead of $v,r$. Double null coordinates are characterized by $g_{UU}=g_{VV}=0$.
$0$ $=g_{VV}=\frac{\partial v}{\partial V}\frac{\partial v}{\partial V}g_{vv}+2\frac{\partial r}{\partial V}\frac{\partial v}{\partial V}g_{rv}$ (II.103)
$=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)+2\frac{\partial r}{\partial V}$ (II.104)
@@
$\frac{\partial r}{\partial V}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (II.105)
@@
$G_{VV}$ $=\frac{\partial v}{\partial V}\frac{\partial v}{\partial V}G_{vv}+2\frac{\partial r}{\partial V}\frac{\partial v}{\partial V}G_{rv}$ (II.106)
$=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)+\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{Q^{2}}{r^{4}}-\frac{Q^{2}}{r^{4}}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (II.107)
$=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)$ (II.108)
@@
$G^{vv}=g^{v\alpha}g^{v\beta}G_{\alpha\beta}=0$ (II.113)
@@
$G^{VV}=\frac{\partial V}{\partial v}\frac{\partial V}{\partial v}G^{vv}=0$ (II.114)
@@
$\widetilde{T}_{uu}=0$ (II.116)
II.8 The source terms in the evolution equations
As mentioned in Eqs. (II.45,II.49), in the absence of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ (namely in double null fluid model) and with constant charge ($Q=const$), $\widetilde{T}_{uu,v}=\widetilde{T}_{vv,u}=0$ – that is, both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are individually conserved. Since our initial values are taken from the Vaidya geometry near the EH, and since $\widetilde{T}_{uu}$ vanishes there (as is evident from regularity conditions), this implies that $\widetilde{T}_{uu}=0$ throughout. We also have $\widetilde{T}_{vv}=2M_{,v}$ , see Eq. (II.109).
@@ II.8 The source terms in the evolution equations
$Z_{R}=8\pi RT_{uv}^{(SC)}$ (II.119)
$Z_{S}=-8\pi\left(\frac{1}{2}T_{uv}^{(SC)}+\frac{L_{0}e^{S}}{R^{\frac{3}{2}}}T_{\theta\theta}^{(SC)}\right)$ (II.120)
@@
Obviously, in order to run a numerical simulation, we have to make a concrete choice of $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$ (as functions of $u$ and $v$). Unfortunately, these two semiclassical functions $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ have not been computed yet inside black holes. Naively one might view this lack of knowledge of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ as a serious obstacle. This is not the case, however, as we now explain:
The purpose of our numerical simulation is to test the validity of an analytical approximation for the evolving semiclassical metric (particularly near the IH), which will be described later (see Sec. IV). This analytical approximation only depends on a set of four main semiclassical input functions: $M(v)$ (closely related to $\widetilde{T}_{vv}^{(+)}(v)$) and $Q^{v}(v)$ (closely related to $J_{v}^{(+)}(v)\propto Q_{,v}^{v}$) at the event horizon – and $\widetilde{T}_{uu}^{(-)}(v),$$\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon, both defined with respect to a background RN metric with parameters $M(v)$,$Q(v)$ . This analytical approximation does not depend on $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$. [The actual values of $\widetilde{T}_{uu}^{(-)},$$\widetilde{T}_{vv}^{(-)}$ are affected by the presence of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ between the two horizon; but nevertheless, once $\widetilde{T}_{uu}^{(-)},\text{$\widetilde{T}_{vv}^{(-)}$}$ are known (and indeed they have been computed, e.g. in Ref. [13]), we no longer need $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ for the mentioned analytical approximation.]
The way we proceed in our numerical investigation is therefore the following: We arbitrarily guess the two semiclassical quantities $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ (as functions of $u,v$), and then numerically evolve the evolution equations (starting from initial data prescribed near the EH). In this process we obtain the evolving unknown functions $R(u,v)$ and $S(u,v)$ – and thereby also the two near-IH flux functions $\widetilde{T}_{uu}^{(-)},$$\widetilde{T}_{vv}^{(-)}$. We can then substitute these latter flux functions in the analytical approximation, and compare it to the actual numerically-evolving functions $R(u,v)$ and $S(u,v)$. This method allows us to numerically test the analytical approximation for any (quite arbitrary) choice of $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$.
III Inputs from semiclassical theory
III.1 Hawking radiation

### v2

@@ 2.5 Ingoing Eddington coordinates
$r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (2.56)
@@
$\frac{\partial r_{*}}{\partial r}=1+\frac{1}{2\kappa_{+}\left(r-r_{+}\right)}-\frac{1}{2\kappa_{-}\left(r-r_{-}\right)}=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}$ (2.57)
@@
$\frac{\partial t}{\partial v}=1,\quad\frac{\partial r}{\partial v}=0$ (2.60)
$\frac{\partial t}{\partial r}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1},\quad\frac{\partial r}{\partial r}=1$ (2.61)
@@
$g_{vv}=\frac{\partial t}{\partial v}\frac{\partial t}{\partial v}g_{tt}+\frac{\partial r}{\partial v}\frac{\partial r}{\partial v}g_{rr}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.62)
$g_{rr}$ $=\frac{\partial t}{\partial r}\frac{\partial t}{\partial r}g_{tt}+\frac{\partial r}{\partial r}\frac{\partial r}{\partial r}g_{rr}$ (2.63)
$=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.64)
$+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}=0$ (2.65)
@@
$g_{rv}$ $=\frac{\partial t}{\partial r}\frac{\partial t}{\partial v}g_{tt}+\frac{\partial r}{\partial r}\frac{\partial r}{\partial v}g_{rr}$ (2.66)
$=\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)=1$ (2.67)
@@
$ds^{2}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)dv^{2}+2dr\;dv+r^{2}d\Omega^{2}$ (2.68)
2.6 RN in double null Eddington coordinates
@@
$\frac{\partial r_{*}}{\partial r}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)^{-1}$ (2.69)
@@
$g_{r_{*}r_{*}}=g_{rr}\frac{\partial r}{\partial r_{*}}\frac{\partial r}{\partial r_{*}}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)$ (2.70)
@@
$ds^{2}=\left(1-\frac{2M}{r(r_{*})}+\frac{Q^{2}}{r^{2}(r_{*})}\right)\left(-dt^{2}+dr_{*}^{2}\right)+r^{2}(r_{*})d\Omega^{2}$ (2.71)
@@
$du\cdot dv=dr_{*}^{2}-dt^{2}$ (2.78)
Substituting this in the metric Eq. (2.71), the Reissner–Nordström geometry in double null coordinates is found to take the form:
$ds^{2}=\left(1-\frac{2M}{r(u,v)}+\frac{Q^{2}}{r^{2}(u,v)}\right)du\;dv+r^{2}(u,v)d\Omega^{2}$ (2.79)
@@
$g_{uv}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.80)
@@
$\frac{u+v}{2}=r_{*}=r-M+\frac{1}{2\kappa_{+}}\log\left|\frac{r-r_{+}}{M-r_{+}}\right|-\frac{1}{2\kappa_{-}}\log\left|\frac{r-r_{-}}{M-r_{-}}\right|$ (2.81)
@@
$\frac{u+v}{2}=r_{*}=const-\frac{1}{2\kappa_{-}}\log\left|r-r_{-}\right|$ (2.82)
@@
$r-r_{-}=const\cdot e^{-2\kappa_{-}r_{*}}=const\cdot e^{-\kappa_{-}\left(u+v\right)}$ (2.83)
The metric component $g_{uv}$ (Eq. (2.80)) around $r=r_{-}$ is (recalling also Eq. (2.1,2.55)):
$2\frac{e^{S}L_{0}}{r}$ $=-2g_{uv}$ (2.84)
$=-\frac{\left(r-r_{-}\right)\left(r-r_{+}\right)}{r^{2}}\approx-\frac{\left(r-r_{-}\right)\left(r_{-}-r_{+}\right)}{r_{-}^{2}}$ (2.85)
$=2\kappa_{-}\left(r-r_{-}\right)$ (2.86)
@@
$const\cdot 2\kappa_{-}e^{-\kappa_{-}\left(u+v\right)}=2\kappa_{-}\left(r-r_{-}\right)\approx 2\frac{e^{s}L_{0}}{r_{-}}$ (2.87)
@@
$S=-\kappa_{-}\left(u+v\right)+const$ (2.88)
@@
$S_{,u}=S_{,v}=-\kappa_{-}$ (2.89)
near the IH. Taking the derivative of Eq. (2.83) with respect to $r_{*},$we get near the IH:
$r_{,r_{*}}=const\cdot e^{-2\kappa_{-}r_{*}}$ (2.90)
@@
$R_{,u}=R_{,r}r_{,r_{*}}r_{*,u}=2r\cdot const\cdot e^{-2\kappa_{-}r_{*}}\frac{1}{2}=const\cdot e^{-2\kappa_{-}r_{*}}\rightarrow 0$ (2.91)
@@
$R_{,v}=R_{,r}r_{,r_{*}}r_{*,v}=const\cdot e^{-2\kappa_{-}r_{*}}\rightarrow 0$ (2.92)
@@
$r_{*}=const+\frac{1}{2\kappa_{+}}\log\left|r-r_{+}\right|$ (2.93)
@@
$r-r_{+}=const\cdot e^{2\kappa_{+}r_{*}}=const\cdot e^{\kappa_{+}\left(u+v\right)}$ (2.94)
@@
$2\frac{e^{S}L_{0}}{r}$ $=-2g_{uv}$ (2.95)
$=-\frac{\left(r-r_{+}\right)\left(r-r_{-}\right)}{r^{2}}\approx-\frac{\left(r-r_{+}\right)\left(r_{+}-r_{-}\right)}{r_{-}^{2}}$ (2.96)
$=-2\kappa_{+}\left(r-r_{+}\right)$ (2.97)
@@
$const\cdot 2\kappa_{+}e^{\kappa_{+}\left(u+v\right)}=2\kappa_{+}\left(r-r_{+}\right)\approx 2g_{uv}$ (2.98)
2.7 Vaidya metric
@@
$ds^{2}=-\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}$ (2.99)
Note that this metric, when taking $M(v)=const$ and $Q(v)=const$, is exactly equal to Eq. (2.68) ingoing Eddington metric.
The $G_{vv}$ component of the Einstein tensor of the Vaidya metric is:
$G_{vv}=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)+\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\frac{Q^{2}}{r^{4}}$ (2.100)
The $G_{rv}$ component of the Einstein tensor of the Vaidya metric is:
$G_{rv}=-\frac{Q^{2}}{r^{4}}$ (2.101)
We wish to know $G_{vv}$ in double null coordinates $(v,u)$, instead of $v,r$. Double null coordinates are characterized by $g_{uu}=g_{vv}=0$.
$0$ $=g_{vv}=\frac{\partial v}{\partial v}\frac{\partial v}{\partial v}g_{vv}+2\frac{\partial r}{\partial v}\frac{\partial v}{\partial v}g_{rv}$ (2.102)
$=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)+2\frac{\partial r}{\partial v}$ (2.103)
@@
$\frac{\partial r}{\partial v}=\frac{1}{2}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (2.104)
@@
$G_{vv}$ $=\frac{\partial v}{\partial v}\frac{\partial v}{\partial v}G_{vv}+2\frac{\partial r}{\partial v}\frac{\partial v}{\partial v}G_{rv}$ (2.105)
$=\frac{1}{r^{2}}\left(2M_{,v}-2\frac{1}{r}QQ_{,v}\right)$ (2.106)
@@
$G^{vv}=g^{v\alpha}g^{v\beta}G_{\alpha\beta}=0$ (2.111)
@@
$G^{vv}=\frac{\partial v}{\partial v}\frac{\partial v}{\partial v}G^{vv}=0$ (2.112)
@@
$\widetilde{T}_{uu}=0$ (2.114)
2.8 The source terms in the evolution equations
As mentioned in Eqs. (2.44,2.48), in the absence of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ (namely in double null fluid model) and with constant charge ($Q=const$), $\widetilde{T}_{uu,v}=\widetilde{T}_{vv,u}=0$ – that is, both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are individually conserved. Since our initial values are taken from the Vaidya geometry near the EH, and since $\widetilde{T}_{uu}$ vanishes there (as is evident from regularity conditions), this implies that $\widetilde{T}_{uu}=0$ throughout. We also have $\widetilde{T}_{vv}=2M_{,v}$ , see Eq. (2.107).
@@ 2.8 The source terms in the evolution equations
$Z_{R}=8\pi RT_{uv}^{(SC)}$ (2.117)
$Z_{S}=-8\pi\left(\frac{1}{2}T_{uv}^{(SC)}+\frac{L_{0}e^{S}}{R^{\frac{3}{2}}}T_{\theta\theta}^{(SC)}\right)$ (2.118)
@@
Obviously, in order to run a numerical simulation, we have to make a concrete choice of $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$ (as functions of $u$ and $v$). Unfortunately, these two semiclassical functions $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ have not been computed yet inside black holes. Naively one might view this lack of knowledge of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ as a crucial obstacle. This is not the case, however, as we now explain:
The purpose of our numerical simulation is to test the validity of an analytical approximation for the evolving semiclassical metric (particularly near the IH), which will be described later (see Sec. 4). This analytical approximation only depends on a set of four main semiclassical input functions: $M(v)$ (closely related to $\widetilde{T}_{vv}^{(+)}(v)$) and $Q^{v}(v)$ (closely related to $J_{v}^{(+)}(v)\propto Q_{,v}^{v}$) at the event horizon – and $\widetilde{T}_{uu}^{(-)}(v),$$\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ at the inner horizon, both defined with respect to a background RN metric with parameters $M(v)$,$Q(v)$ . This analytical approximation does not depend on $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$. [The actual values of $\widetilde{T}_{uu}^{(-)},$$\widetilde{T}_{vv}^{(-)}$ are affected by the presence of $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ between the two horizon; but nevertheless, once $\widetilde{T}_{uu}^{(-)},\text{$\widetilde{T}_{vv}^{(-)}$}$ are known (and indeed they have been computed, e.g. in Ref. [13]), we no longer need $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ for the mentioned analytical approximation.]
The way we proceed in our numerical investigation is therefore the following: We arbitrarily guess the two semiclassical quantities $T_{uv}^{(SC)},T_{\theta\theta}^{(SC)}$ (as functions of $u,v$), and then numerically evolve the evolution equations (starting from initial data prescribed near the EH). The choice of these two RSET components is given in subsection 3.6. (Overall we numerically test 7 different configurations for these RSET components – all of them are of order comparable to the fluxes associated with the evaporation process.) In this process we obtain the evolving unknown functions $R(u,v)$ and $S(u,v)$ – and thereby also the two near-IH flux functions $\widetilde{T}_{uu}^{(-)},$$\widetilde{T}_{vv}^{(-)}$ from the constraint equations Eqs. (2.42, 2.43) . We can then substitute these two flux functions in the analytical approximation, and compare it to the actual numerically-evolving functions $R(u,v)$ and $S(u,v)$. This method allows us to numerically test the analytical approximation for any (quite arbitrary) choice of $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$.
3 Inputs from semiclassical theory
3.1 Hawking radiation